OpenStudy (anonymous):

find the limit as x approaches 1 (2-x)/(x-1)^2 the answer is infinity, but why is it that? can someone explain?

hartnn (hartnn):

simply put x=1 what u get?

OpenStudy (anonymous):

1/0, undefined.

hartnn (hartnn):

thats correct. the limit value = 1/0 = infinity or in other words, the limit is undefined.

OpenStudy (anonymous):

ohh okay :) thanks!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

something else needs to be said, because \(\frac{1}{0}\) is not infinity, even in the limit

OpenStudy (anonymous):

the reason it is \(+\infty\) instead of say \(-\infty\) or neither is because the denominator and numerator are both positive as \(x\) gets close to zero that is because the numerator is close to 1, and the denominator is positive, since it is a square

OpenStudy (anonymous):

I'm confused?

hartnn (hartnn):

did u get that denominator is always positive?

OpenStudy (anonymous):

it needs to be positive or the '0' is positive?

hartnn (hartnn):

since it is a square of a number and a square is always positive thats why... now do u know why numerator is always positive ?

OpenStudy (anonymous):

because x is positive?

hartnn (hartnn):

because x is very near to 1 ..... like 0.999999 or 1.00000001 so 2-x must be positive.. ok ? hence it will be +infinity and not -infinity.

OpenStudy (anonymous):

okay, but the right work to solve the prob is : 2-1/(1-1)^2=1/0=inf?

hartnn (hartnn):

yup, to be more accurate, just add + sign +inf.

OpenStudy (anonymous):

it is positive because if \(x\) is any number not including zero, \((x-1)^2>0\)

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

and don't forget that when you are taking a limit, you are not plugging the number in. that is true no matter what you get. the limit \(\lim_{x\to a}\) automatically assumes that \(x\) is NOT \(a\) but rather is close to \(a\)

OpenStudy (anonymous):

so the denominator is not zero, but close to zero, and it is positive because it is a square. this would not be the case for example for \[\lim_{x\to 1}\frac{x+1}{x-1}\]

OpenStudy (anonymous):

in that case the limit from the right would be positive infinity, but the limit from the left would be negative infinity, so the limit does not exist

OpenStudy (anonymous):

so plugging in 1 for x is the incorrect way to solve it? what is the correct way then?